• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name gilman

    Log entry time 09:49:03 on October 8,2002

    Entry number 87629


    1) We think we understand why the singles rates are so large; it has largely
    been neutrals leading to triggers but no tracks, probably most from the
    radiator which is just at the edge of the acceptence.

    2) We now are running at ~500 Hz T3s on the left arm. There appear to
    be about 50 Hz of VDC travks, and 15 Hz of FPP scatters.

    3) The physics replay should, to be clean, have cuts at Eg>1970 MeV,
    to get rid of any gamma d --> p n pi0 backgrounds, and y_tgt < about 2.5 cm
    or so, to get rid of things from the radiator and target end caps.

    4) There is no way at this point to assess the flow diverter backgrounds.

    Conclusion: it looks like we should be getting good analyzable data.